OpenStudy (anonymous):

Please help me simplify (1+tan^2y)(cos2y-1) Without using sec as the first part because i havent learned that yet. Do i FOIL it? or to I replace cos2y with one of the formulas?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

1+tan^2 y=1+sin^2 y/cos^2 y=(cos^2 y+sin^2 y)/cos^2 y=1/cos^y And cos2y-1=cos^2y-sin^2y-(cos^2y+sin^2y)=-2sin^2y then you can have a good conclusion.